Find how many quarts of 6% butterfat milk and 3% butterfat milk should be mixed to yield 30 quarts of 5% butter milk
Answer by stanbon(75887)

Find how many quarts of 6% butterfat milk and 3% butterfat milk should be mixed to yield 30 quarts of 5% butter milk
-----
Equation:
fat + fat = fat
0.06x + 0.03(30-x) = 0.05*30
----
Multiply thru by 100.
----
6x + 3*30 - 3x = 5*30
---
3x = 2*30
---
x = 20 quarts (amt. of 6% butterfat milk needed)
=====================
Cheers,
Stan H.
=====================